One week 10 copies of the novel The DaVinci Code by Dan Brown were sold for every 1.9 copies of John Grogan's Maryley and Me that were sold. If a total of 3570 copies of the two books were sold, how many copies of each were sold?
Answer by stanbon(75887)

One week 10 copies of the novel The DaVinci Code by Dan Brown were sold for every 1.9 copies of John Grogan's Marley and Me that were sold. If a total of 3570 copies of the two books were sold, how many copies of each were sold?
--------------
10:1.9 is the same as 10x:1.9x where 10 is the number of DaVinci sold.
---
Equation:
10x + 1.9x = 3570
11.9x = 3570
x = 3570/11.9
x = 300
---------------
DaVinci = 10x = 10*300 = 3000
Marly = 1.9x = 570
==============================
Cheers,
Stan H.